OpenStudy (anonymous):

The inverse of the matrix \left[\begin{matrix}3 & 2 \\ 8 & 5\end{matrix}\right] ? The discriminant is 1/-1 , so now what do I do to invert this?

OpenStudy (anonymous):

Shouldn't this be simple? Any suggestions? Just need the inverse of \[\left[\begin{matrix}3 & 2 \\ 8 & 5\end{matrix}\right]\]

OpenStudy (anonymous):

is it \[\left[\begin{matrix}-5 & 2 \\ 8 & -3\end{matrix}\right]\]